OpenStudy (shadydare):

PLEASE HELP!!! Which is a correct comparison of the fractions? A.15/18<25/30 B.15/18>25/30 C.15/18=25/30

OpenStudy (shadydare):

PLEASE HELP!!!

OpenStudy (shadydare):

@NikkiBxox

OpenStudy (nikkibxox):

Alright, Do you know how to simplify fractions? If so, simply 15/18 to its simpilist form and 25/30 the same.

OpenStudy (shadydare):

ok hold on...

OpenStudy (shadydare):

15/18 is 5/6 25/13 is 14/15 I think...

OpenStudy (shadydare):

lol oops **25/30 xD

OpenStudy (nikkibxox):

Try the second one again. The first on you did 15/3 and got 5 correct! 18/3 and got 6 So the GCF (greatest common factor) of 25/30 is 5. Divide 25/5 to get 5 and 30/5 to get 6 so you would then have 15/18= 5/6 and 25/30= 5/6 so your answer is???

OpenStudy (shadydare):

c?

OpenStudy (nikkibxox):

Lamens terms: 15 divided by 3, you get 5. 18 divided by 3, you get 6. Therefor simplified 15/18=5/6 25 divided by 5, you get 5. 30 divided by 5, you get 6. So 25/30= 5/6 Just remember ALWAYS MULTIPLY THE NUMBERATOR AND DENOMINATER BY THE SAME THING! If you remember that you'll do great!! So yes, 5/6=5/6 and 15/18=25/30

OpenStudy (shadydare):

ok thanks :D

OpenStudy (nikkibxox):

Anytime!